Activity in Limit Theorems



Directions: Assume the following.



1. lim f(x) = 3/4


x→c



2. lim g(x) = 12


x→c



3. lim h(x) = -3


x→c






1
Expert's answer
2022-03-14T07:24:24-0400

Given that


"\lim_{x\rightarrow c}f(x)=\frac{3}{4}" "\lim_{x\rightarrow c}g(x)=12" "\lim_{x\rightarrow c}h(x)=-3"


For these values there are no questions, however, a few of the examples are below which will explain how we can use these values using limit theorems.


Ex 1. "\lim_{x\rightarrow c}f(x)+2\lim_{x\rightarrow c}g(x)"


To find "\lim_{x\rightarrow c}f(x)+2\lim_{x\rightarrow c}g(x)" , we use the given values,

"\lim_{x\rightarrow c}f(x)=\frac{3}{4}" and "\lim_{x\rightarrow c}g(x)=12"


Therefore,


"=\frac{3}{4}+2(12)=\frac{99}{4}"




Ex 2. "\lim_{x\rightarrow c}f(x)+2\lim_{x\rightarrow c}g(x)"


To find "\lim_{x\rightarrow c}f(x)+2\lim_{x\rightarrow c}g(x)" , we use the given values,

"\lim_{x\rightarrow c}f(x)=\frac{3}{4}" and "\lim_{x\rightarrow c}g(x)=12"


Therefore,


"=\frac{3}{4}+2(12)=\frac{99}{4}"



Ex 3. "lim_{x\rightarrow c}\frac{f(x)-3g(x)}{h(x)}"


"=\frac{\lim_{x\rightarrow c}f(x)-3\lim_{x\rightarrow c}g(x)}{\lim_{x\rightarrow c}f(x)}"


"=\frac{\frac{3}{4}-(3)(12)}{-3}"


"=\frac{47}{4}"



Ex 4. "lim_{x\rightarrow c}[h(x)-4f(x)]"


"=lim_{x\rightarrow c}h(x)-4lim_{x\rightarrow c}f(x)"

"=(-3)-4(\frac{3}{4})"


"=-3-3"


"=-6"